# EasyText: Controllable Diffusion Transformer for Multilingual Text Rendering

### v1

@@ Abstract
Generating accurate multilingual text with diffusion models has long been desired but remains challenging. Recent methods have made progress in rendering text in a single language, but rendering arbitrary languages is still an unexplored area. This paper introduces EasyText, a text rendering framework based on DiT (Diffusion Transformer), which connects denoising latents with multilingual character tokens encoded as character tokens. We propose character positioning encoding and position encoding interpolation techniques to achieve controllable and precise text rendering. Additionally, we construct a large-scale synthetic text image dataset with 1 million multilingual image-text annotations as well as a high-quality dataset of 20K annotated images, which are used for pretraining and fine-tuning respectively. Extensive experiments and evaluations demonstrate the effectiveness and advancement of our approach in multilingual text rendering, visual quality, and layout-aware text integration. Project page: https://github.com/songyiren725/EasyText
Figure 1: Text-rendered results generated by EasyText, which supports text rendering in over ten languages and produces high-quality results. It can render text either with explicit positional control or in a layout-free manner, and effectively handles curved and slanted regions. The displayed texts do not appear in the prompts; they are included solely to illustrate the intended rendering targets.
1 Introduction
Scene text rendering plays a crucial role in various real-world applications. However, most existing methods such as TextDiffuser [7, 8], Diff-font [14] and modern commercial models like FlUX-dev [20] and Ideogram [1], are primarily limited to English, making multilingual text rendering still a challenging task. GlyphBy-T5-V2 [25] was the first to enable multilingual text rendering by introducing a specially designed glyph encoder and employing a multi-stage training strategy, achieving text rendering in 10 languages on a large-scale dataset of millions of images.
Inspired by how humans learn to write, we gain several key insights: (1) Imitative writing is significantly easier than recalling—humans typically begin by mimicking (i.e., writing with visual reference) before progressing to memory-based writing. (2) Once familiar with one language, humans naturally develop the ability to reproduce text in other unfamiliar languages even without understanding them—treating it more like drawing than writing. Motivated by this, we argue that training AI to “imitate” rather than “recall” is a more efficient and effective strategy for text rendering.
The task of scene text rendering faces several key challenges: (1) Multilingual character modeling is highly complex—for example, Chinese alone has over 30,000 characters. Extending to multilingual settings drastically expands the character space, making joint modeling harder, especially for rare or low-frequency characters with poor rendering accuracy. This is further complicated by language imbalance and font variability. (2) Text-background integration is often unnatural; existing methods struggle to blend rendered text with scene content, leading to visual artifacts such as disconnection or pasted-on effects that undermine image realism. (3) Preserving generative priors is difficult, as fine-tuning on large-scale text-image datasets improves rendering capabilities but often degrades the model’s general image generation ability.
Table 1: Functionality evaluation of EasyText in comparison to other competitors
Functionality Position Control Irregular Regions Multi-Lingual Long Text Rendering Text-Image Blending Unseen Characters Generalization
@@ 1 Introduction
To this end, we present EasyText (shown in Fig. 1), a novel multilingual text image generation framework based on Diffusion Transformers [30, 46]. We encode text into font tokens via a VAE and concatenate them in the latent space with denoised latents. By leveraging the powerful in-context capabilities of DiT, EasyText achieves high-quality and accurate text rendering. Additionally, we propose a simple yet effective position control strategy called Implicit Character Position Alignment, which allows for precise control of character positions through positional encoding interpolation and replacement—enabling both position-aware rendering and layout-free generation.
@@
EasyText is also highly efficient in terms of data usage. Unlike GlyphBy-T5 V1/V2 [24, 25], which rely on contrastive synthetic data or massive real-world scene text datasets, our method simply overlays text randomly on natural images during the pretraining stage to learn glyph features. In this stage, to enable the model to learn glyph imitating rather than simple shape copying, we employ a multi-font mapping approach. Multiple different fonts are overlaid in the synthetic training images, while the condition image uses only a standard font. Afterwards, we fine-tune a lightweight LoRA [16] on only 20K high-quality multilingual scene text images, enhancing the visual consistency between text and background. As shown in Table 1, we highlight some of the capabilities demonstrated by EasyText including: (a) Position control, precisely positioning text in specific locations within an image; (b) Irregular regions, enabling it to handle text rendering in irregular regions such as slanted or curved regions; (c) Multilingual text handling, supporting text generation in multiple languages such as Chinese, English, Japanese, etc; (d) Long text rendering, which can render extended text passages; (e) Text-image blending, which integrates text seamlessly into images, maintaining natural visual consistency; and (f) Unseen character generalization, allowing the model to exhibit generalization capability on unfamiliar and unseen characters.
@@
1.We propose EasyText, a framework that teaches AI to "imitate" rather than "recall," achieving high-quality multilingual text rendering by harnessing the in-context learning power of Diffusion Transformers.
2.We introduce Implicit Character Position Alignment, which precisely controls text placement via lightweight position encoding operations and supports layout-free generation.
3.Extensive experiments demonstrate the effectiveness and simplicity of our method, showing superior performance on challenging scenarios such as long text, multi-text layouts, irregular regions, and unseen characters.
@@ 2.1 Diffusion Models
In recent years, diffusion models [23, 35] have achieved remarkable progress and emerged as a widely used approach for image generation, offering strong capabilities in producing high-quality and diverse visual content through iterative denoising. Their strong capacity for modeling complex data distributions has enabled broad applicability, ranging from image synthesis and editing [4, 15] to video generation [3]. Notable examples include Stable Diffusion [33], its enhanced version SDXL [31], and subsequent variants, which have demonstrated the scalability and effectiveness of text-to-image diffusion models for high-quality image synthesis. Recently, the use of transformer-based denoisers—particularly the Diffusion Transformer (DiT) [30], which replaces U-Net with a transformer backbone—has gained significant traction, and has been adopted in many state-of-the-art models such as FlUX-dev [20], Stable Diffusion 3.5 [11], and PixArt [9].
@@ 2.2 Condition-guided Diffusion Models
Condition-guided diffusion models aim to incorporate external signals—such as spatial structure or semantic reference—into the generative process, enabling control over layout [22, 58], content [55, 56, 50, 57, 59, 37], and identity [49]. These methods facilitate consistent synthesis, alignment with user intent, and support for flexible customization. Early implementations based on U-Net architectures adopted two main paradigms: attention-based conditioning, which integrates semantic features via auxiliary encoders, and residual-based fusion, which injects spatial features into intermediate layers. ControlNet [54] and T2I-Adapter [28] are representative examples that improve spatial consistency and layout control under this framework. With the adoption of transformer-based diffusion models, recent approaches reformulate both semantic and spatial conditions as token sequences. These are integrated into the generation process through multi-modal attention or token concatenation mechanisms, as demonstrated in DiT-based systems like OminiControl [41, tan2025ominicontrol2] . This unified formulation advances the development and application of conditional image generation  [47, 38, 36, 17, 13], improving scalability, simplifying model design, and supporting efficient handling of multiple or high-resolution conditions.
@@ 2.3 Visual Text Generation
Text generation and rendering is a classic task, where early methods mainly relied on generative adversarial networks (GANs) [51, 6, 2, 42] and vector stroke techniques  [43, 39, 40]. Recent work on visual text generation with diffusion models primarily focuses on optimizing text encoding and spatial control mechanisms to improve the fidelity and controllability of rendered text. Character-aware encoders are increasingly adopted: GlyphDraw [27], GlyphControl [52], and AnyText [45] embed glyph or OCR features into conditional inputs, while TextDiffuser-2 [8] leverages character-level tokenization to improve alignment. However, semantic bias and tokenizer limitations still persist. Spatial control has been addressed by introducing explicit layout-related conditions. TextDiffuser [7] and ControlText [18] use segmentation or layout masks, and UDiffText [60] and Brush Your Text [53] refine attention to enforce region-level alignment. For multilingual rendering, works like AnyText [45], GlyphControl [52], and Glyph-ByT5-v2 [25] bypass tokenizers using glyph encoders or tokenizer-free models. Yet, rendering quality remains limited by encoder–diffusion compatibility. Newer models (GlyphDraw2 [26], JoyType [21], AnyText2 [44], FonTS [34], RepText [48]) improve layout precision, while closed models (Kolors 2.0 [19], Seedream 3.0 [12], GPT-4o [29]) show promising results but often suffer from limited spatial controllability and suboptimal performance in multi-text layouts. Despite progress, challenges remain in complex layouts and multilingual text, motivating further integration of glyph, position, and semantic signals.
Figure 2: Overview of EasyText. We adopt a two-stage training strategy: large-scale pretraining for glyph generation and spatial mapping, followed by fine-tuning for visual-text integration and aesthetic refinement. Character positions from the condition input are aligned with target regions via implicit character position alignment, and training proceeds with image-conditioned LoRA.
@@ 3 Method
Section 3.1 introduces FLUX as the base model; Section 3.2 outlines the overall architecture; Section 3.3 describes the target text condition representation; Section 3.4 details the Implicit Character Position Alignment; and Section 3.5 explains the paired dataset construction.
3.1 Preliminary: FLUX
FLUX is an advanced generative framework significantly enhancing generation quality. Departing from traditional U-Net-based diffusion models (such as Stable Diffusion), it adopts a Transformer backbone (DiT) to improve long-sequence modeling and complex semantic understanding. The architecture employs a pre-trained T5 [32] encoder to project textual inputs into dense semantic embeddings $\mathbf{C}_{T}\in\mathbb{R}^{M\times d}$. For visual processing, a variational autoencoder (VAE) performs nonlinear dimensionality reduction on pixel space, generating compact latent representations that preserve essential visual semantics. Additionally, the framework processes encoded noise patches $\mathbf{X}\in\mathbb{R}^{N\times d}$ through rotary position embeddings (RoPE), maintaining spatial coherence during the iterative denoising process. These noise and text tokens are then fused by concatenation to form an image-text token sequence and then fed in parallel into a transformer-based denoising model. The model iteratively denoises and progressively restores a high-quality image aligned with the textual prompt.
@@ 3.2 Overall Architecture
The training pipeline adopts a two-stage strategy. The first stage involves training on a large-scale, synthetically generated dataset with balanced multilingual coverage, enabling the model to learn glyph generation across diverse scripts and accurate spatial mapping. In the second stage, we fine-tune the model on a high-quality annotated dataset to improve both the aesthetic quality of rendered text and its integration with complex scene content. For conditioning, the input text image is encoded into the same latent space as the target image via a VAE. Its positional encoding is aligned with the target region through implicit character position alignment, then concatenated with the target image’s position encoding before being passed to the FLUX DiT block. (Shown in Fig. 2) Image-conditioned LoRA adaptation is then applied, with both the VAE and text encoder kept frozen.
@@ 3.3 Character Representation in EasyFont
First, we establish a unified multilingual representation to addresses the intrinsic typographic differences between writing systems. For alphabetic scripts (e.g., English), we use 64-pixel-high images with widths adaptive to text length, naturally capturing the connected structure of alphabetic word formations. This horizontal layout preserves crucial inter-character relationships and spacing conventions unique to Western typography. Second, for logographic systems (Chinese, Japanese, etc.), we allocate individual 64×64 square images per character. This configuration respects the isolated nature of ideographic characters while maintaining consistent resolution across all glyphs. The square aspect ratio optimally captures the balanced structure inherent to CJK characters, where stroke complexity is uniformly distributed within a square design space.
The image-based character representations are encoded via VAE encoder into latent features. Subsequently, the latent representations undergo patchification through a linear projection layer, which reshapes and projects the spatial features into a sequence of conditioning tokens.
@@ 3.4 Implicit Character Position Alignment
To achieve precise spatial control in font generation, we propose an Implicit Character Position Alignment (ICPA) mechanism through positional encoding extrapolation. Given a conditional image patch $\mathbf{P}_{c}\in\mathbb{R}^{64\times W_{c}\times 3}$ containing source typography features, $W_{c}$ denotes the total width of the conditional image patch, and a target rendering region $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$ in the output image, our method establishes position-aware feature correspondence through positional encoding extrapolation.
Linear Alignment via Affine Transform. Linearly interpolate the position in the condition image (of size $64\times W_{c}$) to the position in the target bounding box $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$. Assuming $u\in[0,W_{c}-1]$ and $v\in[0,63]$ (i.e. the condition image spans indices 00 to $W_{c}$ horizontally and 00 to $63$ vertically), the affine mapping $\mathcal{T}_{\text{aff}}:(u,v)\mapsto(x,y)$ is defined as:
$\mathcal{T}_{\text{aff}}(u,v)=\Big{(}x_{1}+\frac{u}{\,W_{c}-1\,}(x_{2}-x_{1}),\quad y_{1}+\frac{v}{\,63\,}(y_{2}-y_{1})\Big{)}\,,$ (1)
which maps the normalized horizontal coordinate $u/(W_{c}-1)$ to a point between $x_{1}$ and $x_{2}$, and similarly $v/63$ to a point between $y_{1}$ and $y_{2}$, achieving a linear scaling and translation of the patch coordinates into the target domain. The positional encoding is then re-aligned by applying the affine transform to each coordinate in condition image $(u,v)$. Specifically, the transformed positional encoding at the target location $\text{PE}_{t}(x,y)=\mathcal{T}_{\text{aff}}(u,v)$ is assigned from the location in condition image. Thus, the new positional encoding for points in the condition image is given by:
$\text{PE}_{c}(u,v)=\text{PE}_{t}(x,y),$ (2)
where the target position $(x,y)$ is computed via the affine mapping from the patch coordinates.
Nonlinear Alignment via Thin-Plate Spline Interpolation. For nonlinear spatial alignment, we use a Thin-Plate Spline (TPS) interpolation. TPS warping defines a smooth map that exactly interpolates a set of control point correspondences between the source patch and target region, while minimizing bending (second-derivative) distortion. Let $\{(u_{i},v_{i})\}_{i=1}^{K}$ be $K$ landmark points in the source patch $P_{c}$ and $\{(x_{i},y_{i})\}_{i=1}^{K}$ their corresponding positions in the target region $\Omega_{t}$. The TPS mapping $\mathcal{T}_{\text{TPS}}:(u,v)\mapsto(x,y)$ can be expressed as an affine base plus a radial basis deformation term:
$\mathcal{T}_{\text{TPS}}(u,v)=A\begin{pmatrix}u\\ v\\ 1\end{pmatrix}+\sum_{i=1}^{K}w_{i}\,\phi\!\left(\sqrt{(u-u_{i})^{2}+(v-v_{i})^{2}}\right),\quad\phi(r)=r^{2}\log r,$ (3)
@@
$T_{\text{TPS}}(u_{i},v_{i})=(x_{i},y_{i}),i=1,\dots,K,$ (4)
together with additional conditions to ensure a well-posed solution ($\sum_{i=1}^{K}w_{i}=0$, $\sum_{i=1}^{K}w_{i}u_{i}=0$, $\sum_{i=1}^{K}w_{i}v_{i}=0$). This yields a smooth, non-linear warp $\mathcal{T}_{\text{TPS}}$ that maps every source coordinate $(u,v)$ to a target coordinate $(x,y)$, honoring the given correspondences. Similar to linear alignment, target location is obtained via $\text{PE}_{t}(x,y)=\mathcal{T}_{\text{TPS}}(u,v)$. Accordingly, the corresponding positional encoding $\text{PE}_{c}(u,v)$ can be computed as in Equation 2.
Layout-Free Position Alignment via Positional Offset Injection. To enable flexible layout-free rendering, we also introduce an effective positional offset strategy. Instead of aligning coordinates, we shift the positional encoding of the conditional patch by a fixed scalar offset $w_{t}$, which represents the width of the target image. Let $(i,j)$ index the spatial position in $\mathbf{P}_{c}$. Target coordinates $(x^{\prime},y^{\prime})$ are obtained by applying a positional offset to the original coordinates:
$\begin{cases}x^{\prime}=i+w_{t}\\ y^{\prime}=j\end{cases}$ (5)
This ensures positional encoding uniqueness without binding the conditional patch to any specific target location, enabling more flexible text rendering.

### v2

@@ Abstract
Generating accurate multilingual text with diffusion models has long been desired but remains challenging. Recent methods have made progress in rendering text in a single language, but rendering arbitrary languages is still an under-explored area. This paper introduces EasyText, a text rendering framework based on DiT (Diffusion Transformer), which connects denoising latents with multilingual character tokens encoded as character tokens. We propose character positioning encoding and position encoding interpolation techniques to achieve controllable and precise text rendering. Additionally, we construct a large-scale synthetic text image dataset with 1 million multilingual image-text annotations as well as a high-quality dataset of 20K annotated images, which are used for pretraining and fine-tuning respectively. Extensive experiments and evaluations demonstrate the effectiveness and advancement of our approach in multilingual text rendering, visual quality, and layout-aware text integration.
Functionality Position Control Irregular Regions Multi-Lingual Long Text Rendering Text-Image Blending Unseen Characters
@@
Table 1: Functionality evaluation of EasyText in comparison to other competitors.
Codehttps://github.com/songyiren725/EasyText
Datasetshttps://huggingface.co/datasets/lllrrnn/EasyText
Extended versionhttps://arxiv.org/abs/2505.24417
1 Introduction
Scene text rendering is crucial for various real-world applications. However, most existing methods such as TextDiffuser (Chen et al. 2023a, 2024), Diff-font (He et al. 2024) and modern commercial models like FlUX-dev (Labs 2024) and Ideogram (Ideogram Inc. 2025), are primarily limited to English, making multilingual text rendering still a challenging task. Glyph-ByT5-V2 (Liu et al. 2024b) was one of the earliest and most representative works to enable multilingual text rendering by introducing a specially designed glyph encoder and employing a multi-stage training strategy.
Inspired by how humans learn to write, we derive several key insights: (1) Imitative writing is considerably easier than recalling—humans typically begin by mimicking before advancing to memory-based writing. (2) Once familiar with one language, humans naturally develop the ability to reproduce text in other unfamiliar languages even without understanding them—treating it more like drawing than writing. Motivated by this, we argue that training AI to “imitate” rather than “recall” is a more efficient and effective strategy for text rendering.
The task of scene text rendering faces several key challenges: (1) Multilingual character modeling is highly complex—for example, Chinese alone has over 30,000 characters. Extending to multilingual settings drastically expands the character space, making joint modeling harder, especially for rare or low-frequency characters. This is further complicated by language imbalance and font variability. (2) Text-background integration is often unnatural; existing methods struggle to blend rendered text with scene content, leading to visual artifacts such as disconnection or pasted-on effects that undermine image realism. (3) Preserving generative priors is difficult, as fine-tuning on large-scale text-image datasets improves rendering capabilities but often degrades the model’s general image generation ability.
@@ 1 Introduction
To this end, we present EasyText (shown in Fig. 1), a multilingual text image generation framework based on Diffusion Transformers (Peebles and Xie 2023; Vaswani et al. 2017). We encode text into font tokens via a VAE and concatenate them in the latent space with denoised latents. Leveraging the in-context capabilities of DiT, EasyText achieves high-quality and accurate text rendering. Additionally, we propose a simple yet effective position control strategy called Implicit Character Position Alignment, which allows for precise control of character positions through positional encoding interpolation and replacement—enabling both position-aware rendering and layout-free generation.
@@
EasyText is also highly data-efficient. Unlike Glyph-ByT5-V1/V2 (Liu et al. 2024a, b), which rely on contrastive synthetic data or massive real-world text datasets, our method simply overlays text randomly on natural images during the pretraining stage to learn glyph features. To encourage the model to learn glyph imitation rather than simple shape copying, we employ a multi-font mapping approach. Multiple different fonts are overlaid in the synthetic training images, while the condition image uses only a standard font. Afterwards, we fine-tune a lightweight LoRA (Hu et al. 2022) on only 20K high-quality multilingual scene text images, enhancing the visual consistency between text and background. As shown in Table 1, we highlight some of the capabilities demonstrated by EasyText including: (a) Position control, precisely positioning text in specific locations within an image; (b) Irregular regions, enabling it to handle text rendering in irregular regions such as slanted or curved regions; (c) Multilingual text handling, supporting text generation in multiple languages such as Chinese, English, Japanese, etc; (d) Long text rendering, which can render extended text passages; (e) Text-image blending, which integrates text seamlessly into images, maintaining natural visual consistency; and (f) Unseen character generalization, allowing the model to exhibit generalization capability on unfamiliar and unseen characters.
@@
1. We propose EasyText, a framework that teaches AI to “imitate” rather than “recall”, achieving high-quality multilingual text rendering by harnessing the in-context learning power of Diffusion Transformers.
2. We introduce Implicit Character Position Alignment, which precisely controls text placement via position encoding operations and supports layout-free generation.
3. Extensive experiments demonstrate the effectiveness and simplicity of our method, showing superior performance on challenging scenarios such as long text, multi-text layouts, irregular regions, and unseen characters.
Figure 1: Text-rendered results generated by EasyText, which supports text rendering in over ten languages and produces high-quality results. It can render text either with explicit positional control or in a layout-free manner, and effectively handles curved and slanted regions. The additional texts shown below each prompt denote the target texts to be rendered in the image which are not part of the input prompt.
@@ 2.1 Diffusion Models
In recent years, diffusion models (Liang et al. 2024; Song et al. 2020) have achieved remarkable progress and emerged as a widely used approach for image generation, offering strong capabilities in producing high-quality and diverse visual content through iterative denoising. Their strong capacity for modeling complex data distributions has enabled broad applicability, ranging from image synthesis and editing (Brooks et al. 2023; Hertz et al. 2022) to video generation (Bar-Tal et al. 2024). Notable examples include Stable Diffusion (Rombach et al. 2022), its enhanced version SDXL (Podell et al. 2023), and subsequent variants, which have demonstrated the scalability and effectiveness of text-to-image diffusion models for high-quality image synthesis. Recently, the use of transformer-based denoisers—particularly the Diffusion Transformer (DiT) (Peebles and Xie 2023; Feng and Zhang 2023; Feng et al. 2025; Feng and Zhang 2024), which replaces U-Net with a transformer backbone—has gained significant traction, and has been adopted in many state-of-the-art models such as FlUX-dev (Labs 2024), Stable Diffusion 3.5 (Esser et al. 2024), and PixArt (Chen et al. 2023b).
@@ 2.2 Condition-guided Diffusion Models
Condition-guided diffusion models incorporate external signals—such as spatial structure or semantic reference—into the generative process, enabling control over layout (Li et al. 2024b; Zhang et al. 2025), content (Zhang et al. 2024b, c; Wang et al. 2024b; Zhang et al. 2024d, e; Song et al. 2024), motion  (Ma et al. 2025e, 2023, 2024, d, b, c), and identity (Wang et al. 2024a). These methods facilitate consistent synthesis, alignment with user intent, and support for flexible customization. Early implementations based on U-Net architectures adopted two main paradigms: attention-based conditioning, which integrates semantic features via auxiliary encoders, and residual-based fusion, which injects spatial features into intermediate layers. ControlNet (Zhang et al. 2023) and T2I-Adapter (Mou et al. 2024) are representative examples that improve spatial consistency and layout control under this framework. With the adoption of transformer-based diffusion models, recent approaches reformulate both semantic and spatial conditions as token sequences and integrate them into the generation process through multi-modal attention or token concatenation mechanisms, as demonstrated in DiT-based systems like OminiControl (Tan et al. 2024) . This unified formulation advances the development and application of conditional image generation  (Wan et al. 2024; Song et al. 2025b, a; Huang et al. 2025; Guo et al. 2025), improving scalability, simplifying model design, and enabling efficient handling of multiple conditions.
@@ 2.3 Visual Text Generation
Text generation and rendering is a classic task, where early methods mainly relied on generative adversarial networks (GANs) (Wu et al. 2019; Cha et al. 2020; Azadi et al. 2018; Tang et al. 2022) and vector stroke techniques  (Thamizharasan et al. 2024; Song et al. 2023; Song and Zhang 2022). Recent work on visual text generation with diffusion models primarily focuses on optimizing text encoding and spatial control mechanisms to improve the fidelity and controllability of rendered text. Character-aware encoders are increasingly adopted: GlyphDraw (Ma et al. 2023), GlyphControl (Yang et al. 2023), and AnyText (Tuo et al. 2023) embed glyph or OCR features into conditional inputs, while TextDiffuser-2 (Chen et al. 2024) leverages character-level tokenization to improve alignment. However, semantic bias and tokenizer limitations still persist. Spatial control has been addressed by introducing explicit layout-related conditions. TextDiffuser (Chen et al. 2023a) and ControlText (Jiang et al. 2025) use segmentation or layout masks, and UDiffText (Zhao and Lian 2023) and Brush Your Text (Zhang et al. 2024a) refine attention to enforce region-level alignment. For multilingual rendering, works like AnyText (Tuo et al. 2023), GlyphControl (Yang et al. 2023), and Glyph-ByT5-v2 (Liu et al. 2024b) bypass tokenizers using glyph encoders or tokenizer-free models. Yet, rendering quality remains limited by encoder–diffusion compatibility. Newer models (GlyphDraw2 (Ma et al. 2025a), JoyType (Li et al. 2024a), AnyText2 (Tuo et al. 2024), FonTS (Shi et al. 2024), RepText (Wang et al. 2025)) improve layout precision, while closed models (Kolors 2.0 (Kolors Team 2024), WordCon  (Shi et al. 2025), Seedream 3.0 (Gao et al. 2025), GPT-4o (OpenAI 2024)) show promising results but often suffer from limited spatial controllability and suboptimal performance in multi-text layouts. Despite progress, challenges remain in complex layouts and multilingual text, motivating further integration of glyph, position, and semantic signals.
@@ 3 Method
Figure 2: Overview of EasyText. A two-stage training strategy is used: large-scale pretraining for glyph generation and spatial mapping, followed by fine-tuning for visual-text integration and aesthetic refinement. Character positions from the condition input are implicitly aligned with target regions, and training proceeds with image-conditioned LoRA.
The training of our method is based on the open-source base model FLUX (refer to Appendix Section A for details). Section 3.1 outlines the overall architecture; Section 3.2 describes the target text condition representation; Section 3.3 details the Implicit Character Position Alignment; and Section 3.4 explains the paired dataset construction.
@@ 3.1 Overall Architecture
The training pipeline adopts a two-stage strategy. The first stage involves training on a large-scale, synthetically generated dataset with balanced multilingual coverage, enabling the model to learn glyph generation across diverse scripts and accurate spatial mapping. In the second stage, we fine-tune the model on a high-quality annotated dataset to improve both the aesthetic quality of rendered text and its integration with complex scene content. For conditioning, the input condition image ($I_{c}$) is encoded into the same latent space as the target image ($I_{tar}$) via a VAE. Its positional encoding is aligned with the target region through Implicit Character Position Alignment, then concatenated with the target image’s position encoding before being passed to the FLUX DiT block. (Shown in Fig. 2) Image-conditioned LoRA adaptation is then applied, with both the VAE and text encoder kept frozen.
@@ 3.2 Character Representation in EasyFont
First, we establish a unified multilingual representation to address the intrinsic typographic differences between writing systems. For alphabetic scripts (e.g., English), we use 64-pixel-high images with widths adaptive to text length, naturally capturing the connected structure of alphabetic word formations. This horizontal layout preserves crucial inter-character relationships and spacing conventions unique to Western typography. Second, for logographic systems (Chinese, Japanese, etc.), we assign a fixed-size square image to each character. For Chinese characters in particular, the image size is set to 64×64 pixels. This configuration respects the isolated nature of ideographic characters while maintaining consistent resolution. The square aspect ratio optimally captures the balanced structure inherent to characters, where stroke complexity is uniformly distributed within a square design space.
Our condition image efficiently captures the rich visual features of glyphs while providing a significantly more compact representation than many existing methods that use layout inputs matching the target image resolution. By including only the text to be rendered, it typically requires less than one-tenth of the spatial size, substantially reducing computational overhead.
@@ 3.3 Implicit Character Position Alignment
To enable flexible and precise spatial control over the rendered text within the condition image, we introduce an Implicit Character Position Alignment (ICPA) mechanism, which maps the spatial coordinates of the rendered text in the target image (even in irregular regions) onto the corresponding characters in the condition image. Given a conditional image patch $\mathbf{P}_{c}\in\mathbb{R}^{64\times W_{c}\times 3}$ containing source typography features—where $W_{c}$ denotes the total width of the conditional image patch, and a target rendering region $\Omega_{t}$ in the output image, our method establishes position-aware feature correspondence through positional encoding extrapolation.
Linear Alignment via Affine Transform. Linearly interpolate the position in the condition image to the position in the target bounding box (an axis-aligned rectangular region in this case) $\Omega_{t}=[x_{1}:x_{2},y_{1}:y_{2}]$. Assuming $u\in[0,W_{c}-1]$ and $v\in[0,63]$ (i.e. the condition image spans indices 0 to $W_{c}$ horizontally and 0 to $63$ vertically), the affine mapping $\mathcal{T}_{\text{aff}}:(u,v)\mapsto(x,y)$ is defined as:
$\mathcal{T}_{\text{aff}}(u,v)$ $=\Big(x_{1}+\frac{u}{W_{c}-1}(x_{2}-x_{1}),$
$\quad y_{1}+\frac{v}{63}(y_{2}-y_{1})\Big)\,$, (1)
which maps the normalized horizontal coordinate $u/(W_{c}-1)$ to a point between $x_{1}$ and $x_{2}$, and similarly $v/63$ to a point between $y_{1}$ and $y_{2}$, achieving a linear scaling and translation of the coordinates in condition image into the target domain. The positional encoding in condition image is then re-aligned by applying the affine transform, yielding a transformed position $(u^{\prime},v^{\prime})=(x,y)=\mathcal{T}_{\text{aff}}(u,v)$. These transformed coordinates are used to update the positional encoding in condition image, enabling spatial consistency between the condition and target images.
Nonlinear Alignment via Thin-Plate Spline Interpolation. For nonlinear spatial alignment, we adopt Thin-Plate Spline (TPS) interpolation, which defines a smooth mapping that exactly fits a set of control point correspondences between the source patch and the target region, while minimizing second-order bending distortion. Here, $\Omega_{t}$ denotes an irregular region. We obtain $K$ control points along its boundary to form a deformation region that closely fits and fully covers the rendered text. These points act as spatial anchors for the TPS transformation. Let $\{(u_{i},v_{i})\}_{i=1}^{K}$ be $K$ landmark points in the condition image $P_{c}$ and $\{(x_{i},y_{i})\}_{i=1}^{K}$ their corresponding positions in the target region $\Omega_{t}$. The TPS mapping $\mathcal{T}_{\text{TPS}}:(u,v)\mapsto(x,y)$ can be expressed as an affine base plus a radial basis deformation term:
$\mathcal{T}_{\text{TPS}}(u,v)$ $=A\begin{pmatrix}u\\ v\\ 1\end{pmatrix}+\sum_{i=1}^{K}w_{i}\,\phi\!\left(\sqrt{(u-u_{i})^{2}+(v-v_{i})^{2}}\right),$
$\phi(r)$ $=r^{2}\log r,$ (2)
@@
$\mathcal{T}_{\text{TPS}}(u_{i},v_{i})=(x_{i},y_{i}),i=1,\dots,K,$ (3)
together with additional conditions to ensure a well-posed solution ($\sum_{i=1}^{K}w_{i}=0$, $\sum_{i=1}^{K}w_{i}u_{i}=0$, $\sum_{i=1}^{K}w_{i}v_{i}=0$). Similar to linear alignment, transformed position $(u^{\prime},v^{\prime})=(x,y)=\mathcal{T}_{\text{TPS}}(u,v)$ are used to update the positional encoding in condition image.
Layout-Free Position Alignment via Positional Offset Injection. To enable flexible layout-free rendering, we also introduce an effective positional offset strategy. We shift the positional encoding of the conditional image by a fixed scalar offset $w_{t}$, which represents the width of the target image. Let $(u,v)$ index the spatial position in $\mathbf{P}_{c}$. Updated coordinates $(u^{\prime},v^{\prime})$ are obtained by applying a positional offset to the original coordinates:
$\begin{cases}u^{\prime}=u+w_{t}\\ v^{\prime}=v\end{cases}.$ (4)
This ensures positional encoding uniqueness without binding the conditional image to any specific target location, enabling more flexible text rendering.